4x-y=9
x-3y=16
solve graphically

Answer by mananth(16946)

4 x -1 y = 9 .............1
1 x -3 y = 16 .............2
equation1
when x= 0 y= -9 ( 0 , -9 )
when y = 0 x= 2.25 ( 2.25 , 0 )
Equation 2
when x= 0 y= -5.33 ( 0 , -5.33 )
when y = 0 x= 16 ( 16 , 0 )

The solution is (1,-5)
m.ananth@hotmail.ca